# A Fisher-Rao gradient flow for entropic mean-field min-max games

### v1

@@ 2. Main results
Exponential convergence of the single mean-field birth-death flow $(\nu_{t})_{t\geq 0}$ with respect to $\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})$ can be shown to also hold in the setting of [11], however it was not studied in [11], which considered only convergence of $V^{\sigma}(\nu_{t})$ for a convex energy function $V^{\sigma}:\mathcal{P}(\mathbb{R}^{d})\to\mathbb{R}$.
Remark 2.5.
The FR gradient flow can also be interpreted as the continuous-time limit of mirror descent (MD) algorithm with $\operatorname{D_{KL}}$ divergence over the space of probability measures (see, e.g., [11]). In the context of a minimization problem, assuming convexity of the objective function, [2, Theorem $4$] proves that the MD algorithm converges with rate $\mathcal{O}\left(\frac{1}{N}\right)$ to the minimizer, where $N$ is the algorithm’s total number of iterations.
It is crucial to emphasize that the rate $\mathcal{O}\left(\frac{1}{t}\right)$ obtained in Theorem 2.3 is not immediately implicit from the optimization result. Although one might expect that the mirror descent-ascent (MDA) algorithm for a min-max game with convex-concave payoff function converges with rate $\mathcal{O}\left(\frac{1}{N}\right),$ this is not the case because, as stressed in [10], the dynamics of the game are greatly influenced by the order the players move: either simultaneously (players move at the same time) or sequentially (each player moves upon observing the opponents’ moves). The continuous-time analysis does not capture these two situations. Consequently, [10, Theorem $2.6,2.16$] show that the convergence rates of the simultaneous and sequential MDA schemes to mixed Nash equilibria of convex-concave min-max games, measured in the NI error, are $\mathcal{O}\left(N^{-1/2}\right)$ and $\mathcal{O}\left(N^{-2/3}\right),$ respectively.
Remark 2.6.
It is worth commenting that the second result (21) of Theorem 2.3 can be obtained by only requiring $F$ to be convex-concave (see Assumption 1) and setting $\sigma=0$, i.e., the proof of (21) does not use the fact that $V^{\sigma}$ is regularized by $\operatorname{D_{KL}}$ with a positive $\sigma$. Although we are presently unaware of a way to prove the existence of the FR flow (17) for $\sigma=0$ using the Picard iteration technique we employed in Theorem 2.2, we believe that a possible way to prove the existence of the FR flow for $\sigma=0$ could be via a minimizing movement scheme approach in the spirit of [5]. Indeed, the FR gradient flow can be viewed as the zero-step-size limit of a proximal gradient algorithm under the Fisher-Rao distance (see equation $(4.5)$ in [5]). Furthermore, for absolutely continuous measures, one could explicitly write the geodesics minimizing (3) (see Subsection $2.2$ in [5]) along which the iterates generated by the proximal gradient algorithm can be interpolated.
3. Proof of Theorem 2.3
Before we present the proof of Theorem 2.3, we begin with a technical lemma, which extends [11, Lemma 2.5] to the min-max games setup, and which is proved in Section A. The proof of Theorem 2.3 is done in two steps:
@@
  • Second, we show that $\frac{\mathrm{d}}{\mathrm{d}t}\left(\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})\right)$ is bounded from above by $-\frac{\sigma^{2}}{2}\left(\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})\right)$, and then we apply Gronwall’s inequality to obtain exponential convergence. Subsequently, we establish linear convergence for $t\mapsto\operatorname{NI}\left(\frac{1}{t}\int_{0}^{t}\nu_{s}\mathrm{d}s,\frac{1}{t}\int_{0}^{t}\mu_{s}\mathrm{d}s\right)$.
@@
for all $t\geq 0$. Indeed, using (17) and (A.9), we have that
$\Bigg{|}\partial_{t}\left(\nu(x)\log\frac{\nu(x)}{\nu_{t}(x)}\right)\Bigg{|}=\Bigg{|}\nu(x)\frac{\partial_{t}\nu_{t}(x)}{\nu_{t}(x)}\Bigg{|}\\ =\Bigg{|}\nu(x)\left(\frac{\delta F}{\delta\nu}(\nu_{t},\mu_{t},x)+\frac{\sigma^{2}}{2}\log\left(\frac{\nu_{t}(x)}{\pi(x)}\right)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu_{t}|\pi)\right)\Bigg{|}\\ \leq\left(3C_{\nu}+\frac{\sigma^{2}}{2}\left(\max\{|\log r_{1,\nu}|,\log R_{1,\nu}\}+2\log R_{\nu}\right)\right)\nu(x)\coloneqq f(x).$
@@
(22) $\frac{\mathrm{d}}{\mathrm{d}t}\left(\operatorname{D_{KL}}(\nu|\nu_{t})+\operatorname{D_{KL}}(\mu|\mu_{t})\right)\leq V^{\sigma}(\nu,\mu_{t})-V^{\sigma}(\nu_{t},\mu)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu|\nu_{t})-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\mu|\mu_{t}).$
@@
$\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})\leq e^{-\frac{\sigma^{2}}{2}t}\left(\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{0})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{0})\right).$
On the other hand, since $\frac{\sigma^{2}}{2}\left(\operatorname{D_{KL}}(\nu|\nu_{t})+\operatorname{D_{KL}}(\mu|\mu_{t})\right)\geq 0,$ for all $(\nu,\mu),$ (22) becomes
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\operatorname{D_{KL}}(\nu|\nu_{t})+\operatorname{D_{KL}}(\mu|\mu_{t})\right)\leq V^{\sigma}(\nu,\mu_{t})-V^{\sigma}(\nu_{t},\mu).$
Hence, integrating this inequality from 00 to $t>0$ and dividing by $t,$ it follows that
(23) $\frac{1}{t}\int_{0}^{t}\left(V^{\sigma}(\nu_{s},\mu)-V^{\sigma}(\nu,\mu_{s})\right)\mathrm{d}s\leq\frac{1}{t}\left(\operatorname{D_{KL}}(\nu|\nu_{0})+\operatorname{D_{KL}}(\mu|\mu_{0})-\operatorname{D_{KL}}(\nu|\nu_{t})-\operatorname{D_{KL}}(\mu|\mu_{t})\right)\\ \leq\frac{1}{t}\left(\operatorname{D_{KL}}(\nu|\nu_{0})+\operatorname{D_{KL}}(\mu|\mu_{0})\right).$
Then, using the fact that $\nu\mapsto V^{\sigma}(\nu,\mu)$ is convex and $\mu\mapsto V^{\sigma}(\nu,\mu)$ is concave, it follows from Jensen’s inequality that
(24) $\frac{1}{t}\int_{0}^{t}V^{\sigma}(\nu_{s},\mu)\mathrm{d}s-\frac{1}{t}\int_{0}^{t}V^{\sigma}(\nu,\mu_{s})\mathrm{d}s\geq V^{\sigma}\left(\frac{1}{t}\int_{0}^{t}\nu_{s}\mathrm{d}s,\mu\right)-V^{\sigma}\left(\nu,\frac{1}{t}\int_{0}^{t}\mu_{s}\mathrm{d}s\right).$
[…]
$\operatorname{NI}\left(\frac{1}{t}\int_{0}^{t}\nu_{s}\mathrm{d}s,\frac{1}{t}\int_{0}^{t}\mu_{s}\mathrm{d}s\right)\leq\frac{1}{t}\left(\max_{\nu}\operatorname{D_{KL}}(\nu|\nu_{0})+\max_{\mu}\operatorname{D_{KL}}(\mu|\mu_{0})\right).$
@@ Appendix A Technical results and proofs
A.2. Existence and uniqueness of the FR flow
In this subsection, we present the proof of our main result concerning the existence and uniqueness of the Fisher-Rao (FR) flow, i.e., Theorem 2.2. We construct a Picard iteration which is proved to be well-defined in Lemma A.1. Lemma A.2 shows that the Picard iteration is contractive in an appropriate metric. Then in order to conclude the proof of Theorem 2.2 we show the ratio condition (19).
The proof of Theorem 2.2 is an adaptation of the proof of [11, Theorem $2.1$] to the min-max setting (1).
@@ A.2. Existence and uniqueness of the FR flow
(A.1) $\begin{cases}\partial_{t}\nu_{t}(x)=-\left(\frac{\delta F}{\delta\nu}(\nu_{t},\mu_{t},x)+\frac{\sigma^{2}}{2}\log\left(\frac{\nu_{t}(x)}{\pi(x)}\right)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu_{t}|\pi)\right)\nu_{t}(x),\\ \partial_{t}\mu_{t}(y)=\left(\frac{\delta F}{\delta\mu}(\nu_{t},\mu_{t},y)-\frac{\sigma^{2}}{2}\log\left(\frac{\mu_{t}(y)}{\rho(y)}\right)+\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\mu_{t}|\rho)\right)\mu_{t}(y),\end{cases}$
we first notice that (A.1) is equivalent to
(A.2) $\begin{cases}\partial_{t}\log\nu_{t}(x)=-\left(\frac{\delta F}{\delta\nu}(\nu_{t},\mu_{t},x)+\frac{\sigma^{2}}{2}\log\left(\frac{\nu_{t}(x)}{\pi(x)}\right)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu_{t}|\pi)\right),\\ \partial_{t}\log\mu_{t}(y)=\left(\frac{\delta F}{\delta\mu}(\nu_{t},\mu_{t},y)-\frac{\sigma^{2}}{2}\log\left(\frac{\mu_{t}(y)}{\rho(y)}\right)+\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\mu_{t}|\rho)\right).\end{cases}$
By Duhamel’s formula, (A.2) is equivalent to
@@
(A.3) $\begin{split}\log\nu_{t}^{(n)}(x)&=e^{-\frac{\sigma^{2}}{2}t}\log\nu_{0}(x)\\ &-\int_{0}^{t}\frac{\sigma^{2}}{2}e^{-\frac{\sigma^{2}}{2}(t-s)}\left(\frac{2}{\sigma^{2}}\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)-\log\pi(x)-\operatorname{D_{KL}}(\nu_{s}^{(n-1)}|\pi)\right)\mathrm{d}s,\\ \end{split}$
(A.4) $\begin{split}\log\mu_{t}^{(n)}(y)&=e^{-\frac{\sigma^{2}}{2}t}\log\mu_{0}(y)\\ &+\int_{0}^{t}\frac{\sigma^{2}}{2}e^{-\frac{\sigma^{2}}{2}(t-s)}\left(\frac{2}{\sigma^{2}}\frac{\delta F}{\delta\mu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},y)+\log\rho(y)+\operatorname{D_{KL}}(\mu_{s}^{(n-1)}|\rho)\right)\mathrm{d}s.\end{split}$
@@
The sequence of flows $\left((\nu_{t}^{(n)},\mu_{t}^{(n)})_{t\in[0,T]}\right)_{n=0}^{\infty}$ given by (A.3) and (A.4) is well-defined and such that for all $n\geq 1$ and all $t\in[0,T]$ we have
@@
The proof follows from the same induction argument used to prove [11, Lemma $3.1$]. ∎
For fixed $T>0$, we consider the sequence of flows $\left((\nu_{t}^{(n)},\mu_{t}^{(n)})_{t\in[0,T]}\right)_{n=0}^{\infty}$
in $\left(\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]},\mathcal{TV}^{[0,T]}\right),$ where, for any $(\nu_{t},\mu_{t})_{t\in[0,T]}\in\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]}$, the distance $\mathcal{TV}^{[0,T]}$ is defined by
@@
Since $\left(\mathcal{P}(\mathcal{X}),\operatorname{TV}\right)$ is complete, we can apply the argument from [25, Lemma A.5] with $p=1$ to conclude that $\left(\mathcal{P}(\mathcal{X})^{[0,T]},\int_{0}^{T}\operatorname{TV}(\nu_{t},\nu_{t}^{\prime})\mathrm{d}t\right)$ and $\left(\mathcal{P}(\mathcal{Y})^{[0,T]},\int_{0}^{T}\operatorname{TV}(\mu_{t},\mu_{t}^{\prime})\mathrm{d}t\right)$ are complete. Therefore, one can deduce that $\left(\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]},\mathcal{TV}^{[0,T]}\right)$ is also complete. We consider the Picard iteration mapping $\phi\left((\nu_{t}^{(n-1)},\mu_{t}^{(n-1)})_{t\in[0,T]}\right):=(\nu_{t}^{(n)},\mu_{t}^{(n)})_{t\in[0,T]}$ defined via (A.3) and (A.4) and show that $\phi$ is contractive in $\left(\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]},\mathcal{TV}^{[0,T]}\right)$. Then the Banach fixed point theorem will give us the existence and uniqueness of the solution to (17).
@@
The mapping $\phi\left((\nu_{t}^{(n-1)},\mu_{t}^{(n-1)})_{t\in[0,T]}\right):=(\nu_{t}^{(n)},\mu_{t}^{(n)})_{t\in[0,T]}$ defined via (A.3) and (A.4) is contractive in $\left(\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]},\mathcal{TV}^{[0,T]}\right)$.
@@
From (A.3), we have
@@
(A.5) $\operatorname{D_{KL}}(\nu_{t}^{(n)}|\nu_{t}^{(n-1)})=-\int_{0}^{t}\frac{\sigma^{2}}{2}e^{-\frac{\sigma^{2}}{2}(t-s)}\times\\ \times\Bigg{[}\frac{2}{\sigma^{2}}\int_{\mathcal{X}}\left(\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-2)},\mu_{s}^{(n-2)},x)\right)\nu_{t}^{(n)}(\mathrm{d}x)\\ -\operatorname{D_{KL}}(\nu_{s}^{(n-1)}|\pi)+\operatorname{D_{KL}}(\nu_{s}^{(n-2)}|\pi)\Bigg{]}\mathrm{d}s\,.$

### v2

@@ 2 Main results
Exponential convergence of the single mean-field birth-death flow $(\nu_{t})_{t\geq 0}$ with respect to $\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})$ can be shown to also hold in the setting of Liu et al. (2023), however it was not studied in Liu et al. (2023), which considered only convergence of $V^{\sigma}(\nu_{t})$ for a convex energy function $V^{\sigma}:\mathcal{P}(\mathbb{R}^{d})\to\mathbb{R}$.
3 Proof of Theorem 2.3
Before we present the proof of Theorem 2.3, we begin with a technical lemma, which extends Liu et al. (2023, Lemma 2.5) to the min-max games setup, and which is proved in Section A. The proof of Theorem 2.3 is done in two steps:
@@
  • Second, we show that $\frac{\mathrm{d}}{\mathrm{d}t}\left(\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})\right)$ is bounded from above by $-\frac{\sigma^{2}}{2}\left(\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})\right)$, and then we apply Gronwall’s inequality to obtain exponential convergence. Subsequently, we establish exponential convergence for $t\mapsto\operatorname{NI}\left(\nu_{t},\mu_{t}\right).$
@@
for all $t\geq 0$. Indeed, using (19) and (36), we have that
$\Bigg{|}\partial_{t}\left(\nu(x)\log\frac{\nu(x)}{\nu_{t}(x)}\right)\Bigg{|}=\Bigg{|}\nu(x)\frac{\partial_{t}\nu_{t}(x)}{\nu_{t}(x)}\Bigg{|}=\Bigg{|}\nu(x)\left(\frac{\delta F}{\delta\nu}(\nu_{t},\mu_{t},x)+\frac{\sigma^{2}}{2}\log\left(\frac{\nu_{t}(x)}{\pi(x)}\right)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu_{t}|\pi)\right)\Bigg{|}\\ \leq\left(3C_{\nu}+\frac{\sigma^{2}}{2}\left(\max\{|\log r_{1,\nu}|,\log R_{1,\nu}\}+2\log R_{\nu}\right)\right)\nu(x)\coloneqq f(x).$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\operatorname{D_{KL}}(\nu|\nu_{t})+\operatorname{D_{KL}}(\mu|\mu_{t})\right)\leq V^{\sigma}(\nu,\mu_{t})-V^{\sigma}(\nu_{t},\mu)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu|\nu_{t})-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\mu|\mu_{t}).$ (24)
@@
$\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})\leq e^{-\frac{\sigma^{2}}{2}t}\left(\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{0})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{0})\right).$ (25)
To prove convergence in terms of the NI error, first observe that, for any $(\nu,\mu)\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y}),$ we can write
$V^{\sigma}(\nu_{t},\mu)-V^{\sigma}(\nu,\mu_{t})=V^{\sigma}(\nu_{t},\mu)-V^{\sigma}(\nu_{\sigma}^{*},\mu)+V^{\sigma}(\nu_{\sigma}^{*},\mu)-V^{\sigma}(\nu,\mu_{\sigma}^{*})+V^{\sigma}(\nu,\mu_{\sigma}^{*})-V^{\sigma}(\nu,\mu_{t}).$
By Theorem 2.2 and Assumption 2, there exist $C_{1,\sigma},C_{2,\sigma}>0$ depending on $\sigma$ such that
$\left|\frac{\delta V^{\sigma}}{\delta\nu}(\nu_{t},\mu,x)\right|\leq C_{1,\sigma},\quad\left|\frac{\delta V^{\sigma}}{\delta\mu}(\nu,\mu_{t},y)\right|\leq C_{2,\sigma},$
for all $(\nu,\mu)\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y}),$ and all $(x,y)\in\mathcal{X}\times\mathcal{Y}.$ Then, by Lemma 3.1, we have
$V^{\sigma}(\nu_{t},\mu)-V^{\sigma}(\nu_{\sigma}^{*},\mu)$ $\leq\int_{\mathcal{X}}\frac{\delta V^{\sigma}}{\delta\nu}(\nu_{t},\mu,x)(\nu_{t}-\nu_{\sigma}^{*})(\mathrm{d}x)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})$
$\leq C_{1,\sigma}\operatorname{TV}(\nu_{\sigma}^{*},\nu_{t})\leq C_{1,\sigma}\sqrt{2\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})},$
and
$V^{\sigma}(\nu,\mu_{\sigma}^{*})-V^{\sigma}(\nu,\mu_{t})$ $\leq\int_{\mathcal{Y}}\frac{\delta V^{\sigma}}{\delta\mu}(\nu,\mu_{t},y)(\mu_{\sigma}^{*}-\mu_{t})(\mathrm{d}y)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})$
$\leq C_{2,\sigma}\operatorname{TV}(\mu_{\sigma}^{*},\mu_{t})\leq C_{2,\sigma}\sqrt{2\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})},$
where the last inequalities hold due to Pinsker’s inequality and since $\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t}),\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})\geq 0,$ for all $t\geq 0.$ Setting $C_{\sigma}\coloneqq\max\{C_{1,\sigma},C_{2,\sigma}\},$ and using the inequality $\sqrt{a}+\sqrt{b}\leq\sqrt{2(a+b)},$ we get
$V^{\sigma}(\nu_{t},\mu)-V^{\sigma}(\nu,\mu_{t})$ $\leq 2C_{\sigma}\sqrt{\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{t})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{t})}+V^{\sigma}(\nu_{\sigma}^{*},\mu)-V^{\sigma}(\nu,\mu_{\sigma}^{*})$ (26)
$\leq 2C_{\sigma}e^{-\frac{\sigma^{2}}{4}t}\sqrt{\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{0})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{0})}+V^{\sigma}(\nu_{\sigma}^{*},\mu)-V^{\sigma}(\nu,\mu_{\sigma}^{*}),$
where the last inequality follows from (25). Note that since $(\nu_{\sigma}^{*},\mu_{\sigma}^{*})$ is the MNE of (1), we have
$\max_{\mu}V^{\sigma}(\nu_{\sigma}^{*},\mu)=\min_{\nu}V^{\sigma}(\nu,\mu_{\sigma}^{*})=V^{\sigma}(\nu_{\sigma}^{*},\mu_{\sigma}^{*}),$
Therefore, maximizing over $(\nu,\mu)$ in (26) gives
$\operatorname{NI}(\nu_{t},\mu_{t})=\max_{\mu}V^{\sigma}(\nu_{t},\mu)-\min_{\nu}V^{\sigma}(\nu,\mu_{t})\leq 2C_{\sigma}e^{-\frac{\sigma^{2}}{4}t}\sqrt{\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{0})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{0})}.$
Finally, observe that, by (18), we have
$\operatorname{D_{KL}}(\nu_{\sigma}^{*}|\nu_{0})+\operatorname{D_{KL}}(\mu_{\sigma}^{*}|\mu_{0})<\infty.$
@@ Appendix A Technical results and proofs
A.2 Existence and uniqueness of the FR flow
In this subsection, we present the proof of our main result concerning the existence and uniqueness of the Fisher-Rao (FR) flow, i.e., Theorem 2.2. We construct a Picard iteration which is proved to be well-defined in Lemma A.1. Lemma A.2 shows that the Picard iteration mapping admits a unique fixed point in an appropriate metric. Then in order to conclude the proof of Theorem 2.2 we show the ratio condition (21).
The proof of Theorem 2.2 is an adaptation of the proof of Liu et al. (2023, Theorem $2.1$) to the min-max setting (1).
@@ A.2 Existence and uniqueness of the FR flow
$\begin{cases}\partial_{t}\nu_{t}(x)=-\left(\frac{\delta F}{\delta\nu}(\nu_{t},\mu_{t},x)+\frac{\sigma^{2}}{2}\log\left(\frac{\nu_{t}(x)}{\pi(x)}\right)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu_{t}|\pi)\right)\nu_{t}(x),\\ \partial_{t}\mu_{t}(y)=\left(\frac{\delta F}{\delta\mu}(\nu_{t},\mu_{t},y)-\frac{\sigma^{2}}{2}\log\left(\frac{\mu_{t}(y)}{\rho(y)}\right)+\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\mu_{t}|\rho)\right)\mu_{t}(y),\end{cases}$ (27)
we first notice that (27) is equivalent to
$\begin{cases}\partial_{t}\log\nu_{t}(x)=-\left(\frac{\delta F}{\delta\nu}(\nu_{t},\mu_{t},x)+\frac{\sigma^{2}}{2}\log\left(\frac{\nu_{t}(x)}{\pi(x)}\right)-\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\nu_{t}|\pi)\right),\\ \partial_{t}\log\mu_{t}(y)=\left(\frac{\delta F}{\delta\mu}(\nu_{t},\mu_{t},y)-\frac{\sigma^{2}}{2}\log\left(\frac{\mu_{t}(y)}{\rho(y)}\right)+\frac{\sigma^{2}}{2}\operatorname{D_{KL}}(\mu_{t}|\rho)\right).\end{cases}$ (28)
By Duhamel’s formula, (28) is equivalent to
@@
$\begin{split}\log\nu_{t}^{(n)}(x)&=e^{-\frac{\sigma^{2}}{2}t}\log\nu_{0}(x)\\ &-\int_{0}^{t}\frac{\sigma^{2}}{2}e^{-\frac{\sigma^{2}}{2}(t-s)}\left(\frac{2}{\sigma^{2}}\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)-\log\pi(x)-\operatorname{D_{KL}}(\nu_{s}^{(n-1)}|\pi)\right)\mathrm{d}s,\\ \end{split}$ (29)
$\begin{split}\log\mu_{t}^{(n)}(y)&=e^{-\frac{\sigma^{2}}{2}t}\log\mu_{0}(y)\\ &+\int_{0}^{t}\frac{\sigma^{2}}{2}e^{-\frac{\sigma^{2}}{2}(t-s)}\left(\frac{2}{\sigma^{2}}\frac{\delta F}{\delta\mu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},y)+\log\rho(y)+\operatorname{D_{KL}}(\mu_{s}^{(n-1)}|\rho)\right)\mathrm{d}s.\end{split}$ (30)
@@
The sequence of flows $\left((\nu_{t}^{(n)},\mu_{t}^{(n)})_{t\in[0,T]}\right)_{n=0}^{\infty}$ given by (29) and (30) is well-defined and such that for all $n\geq 1$ and all $t\in[0,T]$ we have
@@
The proof follows from the same induction argument used to prove Liu et al. (2023, Lemma $3.1$). ∎
For fixed $T>0$, we consider the sequence of flows $\left((\nu_{t}^{(n)},\mu_{t}^{(n)})_{t\in[0,T]}\right)_{n=0}^{\infty}$ in $\left(\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]},\mathcal{TV}^{[0,T]}\right),$ where, for any $(\nu_{t},\mu_{t})_{t\in[0,T]}\in\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]}$, the distance $\mathcal{TV}^{[0,T]}$ is defined by
@@
Since $\left(\mathcal{P}(\mathcal{X}),\operatorname{TV}\right)$ is complete, we can apply the argument from Šiška & Szpruch (2020, Lemma A.5) with $p=1$ to conclude that $\left(\mathcal{P}(\mathcal{X})^{[0,T]},\int_{0}^{T}\operatorname{TV}(\nu_{t},\nu_{t}^{\prime})\mathrm{d}t\right)$ and $\left(\mathcal{P}(\mathcal{Y})^{[0,T]},\int_{0}^{T}\operatorname{TV}(\mu_{t},\mu_{t}^{\prime})\mathrm{d}t\right)$ are complete. Therefore, one can deduce that $\left(\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]},\mathcal{TV}^{[0,T]}\right)$ is also complete. We consider the Picard iteration mapping $\phi\left((\nu_{t}^{(n-1)},\mu_{t}^{(n-1)})_{t\in[0,T]}\right):=(\nu_{t}^{(n)},\mu_{t}^{(n)})_{t\in[0,T]}$ defined via (29) and (30) and show that $\phi$ admits a unique fixed point $(\nu_{t},\mu_{t})_{t\in[0,T]}$ in $\left(\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]},\mathcal{TV}^{[0,T]}\right),$ which is the solution to (19).
@@
The mapping $\phi$ admits a unique fixed point $(\nu_{t},\mu_{t})_{t\in[0,T]}$ in $\left(\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]},\mathcal{TV}^{[0,T]}\right).$
@@
Step 1: The sequence of flows $\left((\nu_{t}^{(n)},\mu_{t}^{(n)})_{t\in[0,T]}\right)_{n=0}^{\infty}$ is a Cauchy sequence in $\left(\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]},\mathcal{TV}^{[0,T]}\right).$
From (29), we have
@@
$\operatorname{D_{KL}}(\nu_{t}^{(n)}|\nu_{t}^{(n-1)})=-\int_{0}^{t}\frac{\sigma^{2}}{2}e^{-\frac{\sigma^{2}}{2}(t-s)}\Bigg{[}\frac{2}{\sigma^{2}}\int_{\mathcal{X}}\left(\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-2)},\mu_{s}^{(n-2)},x)\right)\nu_{t}^{(n)}(\mathrm{d}x)\\ -\operatorname{D_{KL}}(\nu_{s}^{(n-1)}|\pi)+\operatorname{D_{KL}}(\nu_{s}^{(n-2)}|\pi)\Bigg{]}\mathrm{d}s\,.$ (31)